A 2 m copper wire with square cross-section of 1 mm by 1 mm stretched 1.51 mm when it  supported a 10 kg mass. Suppose you cut a length of this wire 0.2 m long and hang a 10 kg mass  from it. How much will this shore wire stretch?


1
Expert's answer
2021-02-10T01:12:02-0500

First, wed have a wire with long L1 = 2 m and the cross-sectional area is A = 1 mm x 1 mm = 1 x 10-6 m2. This copper wire stretched by s1 = 1.51 mm = 0.00151 m when it supported a weight with mass m = 10 kg. Second, we have the same copper wire with same cross-sectional area, but with long L2 = 0.2 m.

We want to determine the stretch s2. There is a stiffness of the wire itself ks and there is a stiffness of the single interatomic spring ki. The stiffness of the single interatomic spring ki will be the same for the two states at long 2 m and long 0.2 m. So, the strategy is to find ki from ks1 and then use this value to get ks2 and finally use the value of ks2 to find the second stretch s2.

The system here is the mass 10 kg and the surroundings are the Earth (gravitational force) and the wire (tension force). By using the momentum principle we could get the stiffness of this spring when the mass hangs motionless, its momentum is not changing at any time ∆t we find

∆Py = Fnet∆t = 0

As the force in the y-direction, and the tension force is in opposite direction of the gravitational force, therefore, the net force is given by Fnet = F - mg and we could get the next

Fnet∆t = 0

(F - mg)∆t = 0

F - mg = 0

F = mg

kss = mg

ks = (mg) / s

ks1 = (mg) / s1 = (10 kg x 9.8 m/s2) / 0.00151 m = 649 x 102 N/m

The stiffness of the single interatomic bond is very much smaller than the stiffness of the entire wire so we could find the stiffness of the single interatomic bond ki as next

ki = Nbondks / Nchain

Where Nbond is a number of interatomic bonds in one chain and its calculated by

Nbond = Wire's length / Atomic diameter of copper = L1 / d = 2 m / 2.28 x 10-10 m = 8.77 x 109 bonds

And Nchain is the number of chains in the wire and it calculated by

Nchain = Wire's Area / Area of copper atom = Aw / d2 = 1 x 106 m2 / (2.28 x 10-10 m)2 = 1.92 x 1013 chains

Now let us plug our values for Nchain, Nbond and ks1 to get ki

ki = Nbondks1 / Nchain = (8.77 x 109)(649 x 102 N/m) / (1.92 x 1013) = 29.64 N/m

Now we will find our calculations for long L2 = 0.2 m. The number of chains Nchain is the same as cross-sectional area is the same but the number of interatomic bonds Nbond will change as the length is changed

Nbond = Wire's length / Atomic diameter of copper = L2 / d = 0.2 m / 2.28 x 10-10 m = 0.877 x 109 bonds

ks2 = Nchainki / Nbond = (1.92 x 1013)(29.64 N/m) / (0.877 x 109) = 649 x 103 N/m

s2 = (mg) / ks2 = (10 kg x 9.8 m/s2) / (649 x 103 N/m) = 0.151 x 10-3 m = 0.151 mm